OpenStudy (anonymous):

Which is an equation for the given word problem? After Joel deposits $25.50 into his savings account, his balance is $275.00. Which equation can be used to find the amount, n, in his account before he made the deposit? A. 25.50 + n = 275.00 B. 25.50 – n = 275.00 C. n + 275.00 = 25.50 D. n – 275.00 = 25.50

OpenStudy (anonymous):

we know its not A because it would give a negative balance and D would give a even higher balance, C would also give a negative number so its B because you subtract the deposit put it.

OpenStudy (anonymous):

thx

OpenStudy (texaschic101):

I disagree....it is A n = beginning balance n + 25.50 = 275 n = 275 - 25.50 n = 249.50 <--- his starting balance before he added the 25.50

OpenStudy (texaschic101):

@raniaqt